# Energy Window Augmented Plane Waves Approach to Density Functional Theory

### v6

@@ Appendix D Adjusting the basis set size with iteration (review)
Most current DFT calculations for crystalline solids introduce a discrete basis set with the basis set often being indexed by $\mathbf{k},\mathbf{G}$ and for practical reasons is limited to $\mathbf{G}<\mathbf{G}_{max}$, while $\mathbf{k}$ is on a discrete grid. Typically 1000 $\mathbf{k}$ points are taken and roughly 80-100 reciprocal lattice vectors, $\mathbf{G}$ points, per atom [18, 1] are used. In many DFT calculations the dominant contribution to computational time is solving Eq. (19) for a total of $\mathcal{N}$ times till convergence with each run taking $\sim\mathbf{G}_{max}^{9}$ time for a total of $\sim\mathcal{N}\mathbf{G}_{max}^{9}$ computations. Here $\mathcal{N\sim}20$ for many situations [18]. Now at each step from numerical simulations it is known that the error from an exact solution the KS equations is given by [1, 18, 45, 9]:

### v7

@@ Appendix D Adjusting the basis set size with iteration (review)
Most current DFT calculations for crystalline solids introduce a discrete basis set with the basis set often being indexed by $\mathbf{k},\mathbf{G}$ and for practical reasons is limited to $\mathbf{G}<\mathbf{G}_{max}$, while $\mathbf{k}$ is on a discrete grid. Typically several hundred $\mathbf{k}$ points are taken and roughly 80-100 reciprocal lattice vectors, $\mathbf{G}$ points, per atom [19, 1] are used for the case of large MT spheres. In many DFT calculations the dominant contribution to computational time is solving Eq. (19) for a total of $\mathcal{N}$ times till convergence with each run taking $\sim\mathbf{G}_{max}^{9}$ time for a total of $\sim\mathcal{N}\mathbf{G}_{max}^{9}$ computations. Here $\mathcal{N\sim}20$ for many situations [19]. Now at each step from numerical simulations it is known that the error from an exact solution the KS equations is given by [1, 19, 52, 9]: